OpenStudy (anonymous):

show that 12 sin^2x cos^2x = 3/2(1-cos4x)

OpenStudy (anonymous):

2sin^2x = 1-cos2x 2cos^2x = 1+cos2x use the above formulas to simplify

OpenStudy (anonymous):

12 sin^2x cos^2x = 3*2sin^2x *2cos^2x =3*(1-cos2x)(1+cos2x) =3*(1-cos^2 2x) =3*sin^2 2x =3/2 * 2*sin^2 2x =3/2 *(1-cos4x)